# Square Functions Controlling Smoothness and Higher-Order Quantitative Rectifiability

### v1

@@ 3. $I_{\alpha}(BMO)$ Implies Square Function Bounds
(3.13) $\int_{0}^{R}\int_{B_{R}(z)}\left(\overline{\nu}_{1}^{g}(x,r)\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\lesssim\int_{\mathbb{R}^{d}}\left(\hat{g}(\xi)|\xi|\right)^{2}\,\mathrm{d}\xi\approx\int_{\mathbb{R}^{d}}|\nabla g(x)|^{2}\,\mathrm{d}x\lesssim\int|I(x)|^{2}\,\mathrm{d}x+\int|II(x)|^{2}\,\mathrm{d}x$
@@
$|\nabla g(x)|\\ \leq\left|R^{-1}(\nabla\phi)\left(\frac{x-z}{R}\right)\big{(}f(x)-f(z)-\nabla(f\ast\varphi_{R})(z)\cdot(x-z)\big{)}\right|+\left|\varphi\left(\frac{x-z}{R}\right)\big{(}\nabla f(x)-\nabla(f\ast\varphi_{R})(z)\big{)}\right|\\ =:I(x)+II(x)$
@@
$+\int_{0}^{1}R^{d}\left|\nabla(f\ast\varphi_{R})(z)-\nabla(f\ast\varphi_{tR})(z)\right|^{2}\,\mathrm{d}t=T_{1}+T_{2}$
@@
$\left|\nabla(f\ast\varphi_{R})(z)-\nabla(f\ast\varphi_{tR})(z)\right|\\ \left|\int(\nabla f)(z-y)\big{(}\varphi_{R}(y)-\varphi_{tR}(y)\big{)}\,\mathrm{d}y\right|=\left|\int(\nabla f)(z-(Rt)y)\big{(}t^{d}\varphi(ty)-\varphi(y)\big{)}\,\mathrm{d}y\right|\\ \leq\|\nabla f\|_{*}\|t^{d}\varphi(t\cdot)-\varphi(\cdot)\|_{H^{1}}\lesssim\|\nabla f\|_{*}$
@@
Now we consider the cases when $\alpha\neq 1$. We will use Propositions 3.1 and 3.4 along with an approximation argument.
@@
$g(x):=\phi\bigg{(}\frac{x-z}{R}\bigg{)}\big{(}f(x)-f(z)-\nabla f(z)\cdot(x-z)\big{)}$
@@
$\int_{0}^{R}\int_{B_{R}(z)}\left(\frac{\overline{\nu}_{1}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}=\int_{0}^{R}\int_{B_{R}(z)}\left(\frac{\overline{\nu}_{1}^{g}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}$
@@
$f^{t}(x)$ $=(f\ast\psi_{t})(x)$
$g^{t}(x)$ $=\psi\left(\frac{x-z}{R}\right)\big{(}f^{t}(x)-f^{t}(z)-\nabla f^{t}(z)\cdot(x-z)\big{)}$
@@
(3.14) $\int_{\mathbb{R}^{d}}\big{|}|\xi|^{\alpha}\hat{g^{t}}(\xi)\big{|}^{2}\,\mathrm{d}\xi\lesssim\|D_{\alpha}f^{t}\|_{*}^{2}R^{d}$
@@
(3.15) $\int_{0}^{R}\int_{B_{R}(z)}\left(\frac{\overline{\nu}^{f^{t}}_{1}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\lesssim\int_{\mathbb{R}^{d}}\big{(}\hat{g}^{t}(\xi)|\xi|^{\alpha}\big{)}^{2}\,\mathrm{d}\xi$
@@
(3.16) $|\overline{\nu}_{1}^{f^{t}-f}(x,r)^{2}|\lesssim\fint_{B_{r}(x)}\left(\frac{|(f^{t}(y)-f(y))-(f^{t}(x)-f(x))-\nabla(f^{t}(x)-f(x))\cdot(y-x)|}{r}\right)^{2}\,\mathrm{d}y\lesssim t^{2(\alpha-1)}$
@@
$\int_{\sqrt{t}}^{R}\left(\frac{\overline{\nu}_{1}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\lesssim\int_{\sqrt{t}}^{R}\left(\frac{\overline{\nu}_{1}^{f^{t}}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}+R^{d}\int_{\sqrt{t}}^{R}\frac{t^{2\alpha-2}}{r^{2\alpha-1}}\,\mathrm{d}x\,\mathrm{d}r\\ \lesssim\int_{\mathbb{R}^{d}}\big{(}\widehat{g^{t}}(\xi)|\xi|^{\alpha}\big{)}^{2}\,\mathrm{d}\xi+R^{d}t^{\alpha-1}\lesssim\|D_{\alpha}f\|_{*}^{2}R^{d}+R^{d}t^{\alpha-1}$
@@
$f^{t}(x)$ $=(f\ast\psi_{t})(x)$
$g^{t}(x)$ $=\psi\left(\frac{x-z}{R}\right)\big{(}f_{t}(x)-f_{t}(z)\big{)}$
@@
$|\overline{\nu}_{1}^{f^{t}-f}(x,r)^{2}|=\fint_{B_{r}(x)}\left(\frac{|(f^{t}(y)-f(y))-(f^{t}(x)-f(x))|}{r}\right)^{2}\,\mathrm{d}y\leq\frac{t^{2\alpha}}{r^{2}}$
@@
$\int_{\sqrt{t}}^{R}\int_{B_{R}(z)}\left(\frac{\nu_{1}^{f^{t}-f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\lesssim R^{d}t^{\alpha}$
@@
(4.2) $\sup_{Q}\left(\frac{1}{|Q|}\int_{Q}\int_{Q}\frac{|f(x+y)-f(x)|^{2}}{|x-y|^{d+2\alpha}}\,\mathrm{d}y\,\mathrm{d}x\right)^{\frac{1}{2}}=B<\infty$
@@
(4.3) $\int\frac{|f(x)|}{1+|x|^{d+\epsilon}}\,\mathrm{d}x<\infty$
@@
(4.4) $\sup_{Q}\left(\frac{1}{|Q|}\int_{Q}\int_{Q}\frac{|2f(x)-f(x+y)-f(x-y)|^{2}}{|x-y|^{d+2\alpha}}\,\mathrm{d}y\,\mathrm{d}x\right)^{\frac{1}{2}}=B<\infty$
@@
(4.7) $\tilde{\nu}_{1}^{f}(x,r)=\left(\fint_{\frac{r}{2}\leq|y|\leq r}\left(\frac{|f(x+y)-f(x)-(\phi_{r}\ast\nabla f)(x)\cdot y|}{r}\right)^{2}\,\mathrm{d}y\right)^{\frac{1}{2}}$
@@
$\int_{0}^{R}\int_{B_{R}(z)}\left(\frac{\tilde{\nu}_{0}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}x\\ \lesssim\int_{0}^{R}\int_{B_{R}(z)}\fint_{\frac{r}{2}\leq|y|\leq r}\left(\frac{|f(x)-c_{(x,r)}|}{r}\right)^{2}\,\mathrm{d}yr^{1-2\alpha}\,\mathrm{d}x\,\mathrm{d}r\\ +\int_{0}^{R}\int_{B_{R}(z)}\fint_{\frac{r}{2}\leq|y|\leq r}\left(\frac{|f(x+y)-c_{(x,r)}|}{r}\right)^{2}\,\mathrm{d}yr^{1-2\alpha}\,\mathrm{d}x\,\mathrm{d}r\\ =I+II$
@@
$\int_{0}^{R}\int_{B_{R}(z)}\left(\frac{\nu_{0}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\leq C_{1}R^{d}$
@@
[…]
@@
Now we prove the case when $21\leq\alpha<$ 2. We now let $l_{(x,r)}$ denote the linear function minimizing $\nu_{1}^{f}(x,r)$. We note that for any linear function $L(x)$ one has $L(x+y)=L(x)+y\cdot\big{(}\phi_{r}\ast\nabla L(x)\big{)}$. We apply this fact to $l_{(x,r)}$ to obtain the estimate
$\int_{0}^{R}\int_{B_{R}(z)}\left(\frac{\tilde{\nu}_{1}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\lesssim\int_{0}^{R}\int_{B_{R}(z)}\fint_{r/2\leq|y|\leq r}\left(\frac{|f(x+y)-l_{(x,r)}(x+y)|}{r}\right)^{2}\,\mathrm{d}yr^{1-2\alpha}\,\mathrm{d}x\,\mathrm{d}r\\ +\int_{0}^{R}\int_{B_{R}(z)}\fint_{r/2\leq|y|\leq r}\left(\frac{|f(x)-l_{(x,r)}(x)|}{r}\right)^{2}\,\mathrm{d}yr^{1-2\alpha}\,\mathrm{d}x\,\mathrm{d}r\\ +\int_{0}^{R}\int_{B_{R}(z)}\fint_{r/2\leq|y|\leq r}\left(\frac{|y\cdot(\phi_{r}\ast\nabla(f-l_{(x,r)}))(x)|}{r}\right)^{2}\,\mathrm{d}yr^{1-2\alpha}\,\mathrm{d}x\,\mathrm{d}r\\ =I+II+III$
@@
$I\lesssim\int_{0}^{R}\int_{B^{\prime}_{R}(z)}\left(\frac{\nu_{1}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\lesssim C_{1}R^{d}$
@@
$\left(\frac{y\cdot\frac{1}{r}((\nabla\phi_{r})\ast(f-l_{(x,r)})(x)}{r}\right)^{2}\lesssim\left(\fint_{B_{r}(x)}\frac{|f(\zeta)-l_{(x,r)}(\zeta)|}{r}\,\mathrm{d}\zeta\right)^{2}\leq\nu_{1}^{f}(x,r)^{2}$
@@
$\int_{0}^{R}\int_{B^{\prime}_{R}(z)}\left(\frac{\nu^{f}_{\lfloor\alpha\rfloor}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\leq C_{1}R^{d}$
for all $R>0$ and all $z\in\mathbb{R}^{d}$. Additionally, in the case that $21\leq\alpha<$ 2, assume $f$ satisfies 1.15 for all $\epsilon>\alpha$. Then $f\in I_{\alpha}(BMO)$, and $\|D_{\alpha}f\|_{BMO}^{2}\lesssim C_{1}$.
@@
$\int_{0}^{R}\int_{B_{R}(z)}\left(\frac{\tilde{\nu}_{\lfloor\alpha\rfloor}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\lesssim R^{d}$

### v2

@@ 3. $I_{\alpha}(BMO)$ Implies Square Function Bounds
(3.14) $\int_{0}^{R}\int_{B_{R}(z)}\left(\overline{\nu}_{1}^{g}(x,r)\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\lesssim\int_{\mathbb{R}^{d}}\left(\hat{g}(\xi)|\xi|\right)^{2}\,\mathrm{d}\xi\approx\int_{\mathbb{R}^{d}}|\nabla g(x)|^{2}\,\mathrm{d}x\lesssim\int|I(x)|^{2}\,\mathrm{d}x+\int|II(x)|^{2}\,\mathrm{d}x.$
@@
$|\nabla g(x)|\\ \leq\left|R^{-1}(\nabla\phi)\left(\frac{x-z}{R}\right)\big{(}f(x)-f(z)-\nabla(f\ast\varphi_{R})(z)\cdot(x-z)\big{)}\right|+\left|\varphi\left(\frac{x-z}{R}\right)\big{(}\nabla f(x)-\nabla(f\ast\varphi_{R})(z)\big{)}\right|\\ =:I(x)+II(x).$
@@
$+\int_{0}^{1}R^{d}\left|\nabla(f\ast\varphi_{R})(z)-\nabla(f\ast\varphi_{tR})(z)\right|^{2}\,\mathrm{d}t=T_{1}+T_{2}.$
@@
$\left|\nabla(f\ast\varphi_{R})(z)-\nabla(f\ast\varphi_{tR})(z)\right|\\ \left|\int(\nabla f)(z-y)\big{(}\varphi_{R}(y)-\varphi_{tR}(y)\big{)}\,\mathrm{d}y\right|=\left|\int(\nabla f)(z-(Rt)y)\big{(}t^{d}\varphi(ty)-\varphi(y)\big{)}\,\mathrm{d}y\right|\\ \leq\|\nabla f\|_{*}\|t^{d}\varphi(t\cdot)-\varphi(\cdot)\|_{H^{1}}\lesssim\|\nabla f\|_{*}.$
@@
Now we consider the cases when $\alpha\neq 1$. We will use Propositions 3.1 and 3.5 along with an approximation argument.
@@
$g(x):=\phi\bigg{(}\frac{x-z}{R}\bigg{)}\big{(}f(x)-f(z)-\nabla f(z)\cdot(x-z)\big{)}.$
@@
$\int_{0}^{R}\int_{B_{R}(z)}\left(\frac{\overline{\nu}_{1}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}=\int_{0}^{R}\int_{B_{R}(z)}\left(\frac{\overline{\nu}_{1}^{g}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}.$
@@
$f^{t}(x)$ $=(f\ast\psi_{t})(x),$
$g^{t}(x)$ $=\psi\left(\frac{x-z}{R}\right)\big{(}f^{t}(x)-f^{t}(z)-\nabla f^{t}(z)\cdot(x-z)\big{)}.$
@@
(3.15) $\int_{\mathbb{R}^{d}}\big{|}|\xi|^{\alpha}\hat{g^{t}}(\xi)\big{|}^{2}\,\mathrm{d}\xi\lesssim\|D_{\alpha}f^{t}\|_{*}^{2}R^{d}.$
@@
(3.16) $\int_{0}^{R}\int_{B_{R}(z)}\left(\frac{\overline{\nu}^{f^{t}}_{1}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\lesssim\int_{\mathbb{R}^{d}}\big{(}\hat{g}^{t}(\xi)|\xi|^{\alpha}\big{)}^{2}\,\mathrm{d}\xi.$
@@
(3.17) $|\overline{\nu}_{1}^{f^{t}-f}(x,r)^{2}|\lesssim\fint_{B_{r}(x)}\left(\frac{|(f^{t}(y)-f(y))-(f^{t}(x)-f(x))-\nabla(f^{t}(x)-f(x))\cdot(y-x)|}{r}\right)^{2}\,\mathrm{d}y\lesssim t^{2(\alpha-1)}.$
@@
$\int_{\sqrt{t}}^{R}\int_{B_{R}(z)}\left(\frac{\overline{\nu}_{1}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\lesssim\int_{\sqrt{t}}^{R}\int_{B_{R}(z)}\left(\frac{\overline{\nu}_{1}^{f^{t}}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}+R^{d}t^{\alpha-1}\\ \lesssim\int_{\mathbb{R}^{d}}\big{(}\widehat{g^{t}}(\xi)|\xi|^{\alpha}\big{)}^{2}\,\mathrm{d}\xi+R^{d}t^{\alpha-1}\lesssim\|D_{\alpha}f\|_{*}^{2}R^{d}+R^{d}t^{\alpha-1}.$
@@
$f^{t}(x)$ $=(f\ast\psi_{t})(x),$
$g^{t}(x)$ $=\psi\left(\frac{x-z}{R}\right)\big{(}f_{t}(x)-f_{t}(z)\big{)}.$
@@
$|\overline{\nu}_{1}^{f^{t}-f}(x,r)^{2}|=\fint_{B_{r}(x)}\left(\frac{|(f^{t}(y)-f(y))-(f^{t}(x)-f(x))|}{r}\right)^{2}\,\mathrm{d}y\lesssim t^{2\alpha}.$
@@
$\int_{\sqrt{t}}^{R}\int_{B_{R}(z)}\left(\frac{\nu_{1}^{f^{t}-f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\lesssim R^{d}t^{\alpha}.$
@@
(4.2) $\sup_{Q}\left(\frac{1}{|Q|}\int_{Q}\int_{Q}\frac{|f(x+y)-f(x)|^{2}}{|x-y|^{d+2\alpha}}\,\mathrm{d}y\,\mathrm{d}x\right)^{\frac{1}{2}}=B<\infty.$
@@
(4.3) $\int\frac{|f(x)|}{1+|x|^{d+\epsilon}}\,\mathrm{d}x<\infty,$
@@
(4.4) $\sup_{Q}\left(\frac{1}{|Q|}\int_{Q}\int_{Q}\frac{|2f(x)-f(x+y)-f(x-y)|^{2}}{|x-y|^{d+2\alpha}}\,\mathrm{d}y\,\mathrm{d}x\right)^{\frac{1}{2}}=B<\infty.$
@@
(4.7) $\tilde{\nu}_{1}^{f}(x,r)=\left(\fint_{\frac{r}{2}\leq|y|\leq r}\left(\frac{|f(x+y)-f(x)-(\phi_{r}\ast\nabla f)(x)\cdot y|}{r}\right)^{2}\,\mathrm{d}y\right)^{\frac{1}{2}},$
@@
$\int_{0}^{R}\int_{B_{R}(z)}\left(\frac{\tilde{\nu}_{\lfloor\alpha\rfloor}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\lesssim C_{1}R^{d}.$
@@
$\int_{0}^{R}\int_{B_{R}(z)}\left(\frac{\tilde{\nu}_{0}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\\ \lesssim\int_{0}^{R}\int_{B_{R}(z)}\fint_{\frac{r}{2}\leq|y|\leq r}\left(\frac{|f(x)-c_{(x,r)}|}{r}\right)^{2}\,\mathrm{d}yr^{1-2\alpha}\,\mathrm{d}x\,\mathrm{d}r\\ +\int_{0}^{R}\int_{B_{R}(z)}\fint_{\frac{r}{2}\leq|y|\leq r}\left(\frac{|f(x+y)-c_{(x,r)}|}{r}\right)^{2}\,\mathrm{d}yr^{1-2\alpha}\,\mathrm{d}x\,\mathrm{d}r\\ =I+II.$
@@
$\int_{0}^{R}\int_{B_{R}(z)}\left(\frac{\nu_{0}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\leq C_{1}R^{d}.$
@@
$\int_{0}^{R}\int_{B_{2R}(z)}\fint_{\frac{r}{2}\leq|x-y|\leq r}\left(\frac{|f(x)-c_{(x,r)}|}{r}\right)^{2}\,\mathrm{d}xr^{1-2\alpha}\,\mathrm{d}y\,\mathrm{d}r.$
@@
Now we prove the case when $21\leq\alpha<$ 2. We now let $l_{(x,r)}$ denote the affine function minimizing $\nu_{1}^{f}(x,r)$. We note that for any affine function $L(x)$ one has $L(x+y)=L(x)+y\cdot\big{(}\phi_{r}\ast\nabla L(x)\big{)}$. We apply this fact to $l_{(x,r)}$ to obtain the estimate
$\int_{0}^{R}\int_{B_{R}(z)}\left(\frac{\tilde{\nu}_{1}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\lesssim\int_{0}^{R}\int_{B_{R}(z)}\fint_{r/2\leq|y|\leq r}\left(\frac{|f(x+y)-l_{(x,r)}(x+y)|}{r}\right)^{2}\,\mathrm{d}yr^{1-2\alpha}\,\mathrm{d}x\,\mathrm{d}r\\ +\int_{0}^{R}\int_{B_{R}(z)}\fint_{r/2\leq|y|\leq r}\left(\frac{|f(x)-l_{(x,r)}(x)|}{r}\right)^{2}\,\mathrm{d}yr^{1-2\alpha}\,\mathrm{d}x\,\mathrm{d}r\\ +\int_{0}^{R}\int_{B_{R}(z)}\fint_{r/2\leq|y|\leq r}\left(\frac{|y\cdot(\phi_{r}\ast\nabla(f-l_{(x,r)}))(x)|}{r}\right)^{2}\,\mathrm{d}yr^{1-2\alpha}\,\mathrm{d}x\,\mathrm{d}r\\ =I+II+III.$
@@
$I\lesssim\int_{0}^{R}\int_{B^{\prime}_{R}(z)}\left(\frac{\nu_{1}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\lesssim C_{1}R^{d}.$
@@
$\left(\frac{y\cdot\frac{1}{r}((\nabla\phi_{r})\ast(f-l_{(x,r)})(x)}{r}\right)^{2}\lesssim\left(\fint_{B_{r}(x)}\frac{|f(\zeta)-l_{(x,r)}(\zeta)|}{r}\,\mathrm{d}\zeta\right)^{2}\leq\nu_{1}^{f}(x,r)^{2}.$
@@
$\int_{0}^{R}\int_{B^{\prime}_{R}(z)}\left(\frac{\nu^{f}_{\lfloor\alpha\rfloor}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\leq C_{1}R^{d}.$
for all $R>0$ and all $z\in\mathbb{R}^{d}$. Additionally, in the case that $21\leq\alpha<$ 2, assume $f$ satisfies 1.15 for all $\epsilon>\alpha$. Then $f\in I_{\alpha}(BMO)$, and $\|D_{\alpha}f\|_{*}^{2}\lesssim C_{1}$.